# InterIntent: Investigating Social Intelligence of LLMs via Intention Understanding in an Interactive Game Context

### v1

@@ Abstract
Large language models (LLMs) have demonstrated the potential to mimic human social intelligence. However, most studies focus on simplistic and static self-report or performance-based tests, which limits the depth and validity of the analysis. In this paper, we developed a novel framework, InterIntent, to assess LLMs’ social intelligence by mapping their ability to understand and manage intentions in a game setting. We focus on four dimensions of social intelligence: situational awareness, self-regulation, self-awareness, and theory of mind. Each dimension is linked to a specific game task: intention selection, intention following, intention summarization, and intention guessing. Our findings indicate that while LLMs exhibit high proficiency in selecting intentions, achieving an accuracy of 88%, their ability to infer the intentions of others is significantly weaker, trailing human performance by 20%. Additionally, game performance correlates with intention understanding, highlighting the importance of the four components towards success in this game. These findings underline the crucial role of intention understanding in evaluating LLMs’ social intelligence and highlight the potential of using social deduction games as a complex testbed to enhance LLM evaluation. InterIntent contributes a structured approach to bridging the evaluation gap in social intelligence within multiplayer games.
@@ 1 Introduction
Despite adaptations for interactive social environments Zhou et al. (2023c, a); Park et al. (2023), most existing evaluations on LLMs Le et al. (2019); Shapira et al. (2023) focus on straightforward and static daily scenarios without explicitly defined goals for them. To enhance the contexts with complexity and dynamics with clear goals for LLMs, we turn our attention to social deduction games. These games require social interactions among players, providing a more diverse test bed for LLM evaluation Qiao et al. (2023). Existing studies have successfully utilized various games to analyze LLMs’ social behaviors, such as deception and confrontation Liang et al. (2023); Ibraheem et al. (2022); Mansbach et al. (2021); Meta et al. (2022); O’Gara (2023); Xu et al. (2023a, b); Wu et al. (2024). However, these studies often focus on ad-hoc post-analysis of game performance and overlook systematic measurement of social intelligence, limiting the comprehensive understanding of LLMs’ capabilities in social environments.
In this paper, we leverage one of the representatives in social deduction games, Avalon Light et al. (2023a); Wang et al. (2023), as the context for LLM evaluation. Avalon is a social role-playing game based on deduction that heavily relies on conversation, making it an excellent test bed due to its goal-driven objectives and complex mechanisms, which facilitates social interactions among LLMs within a detailed, grounded environment. We systematically evaluate LLMs’ social intelligence across the four aforementioned dimensions. Specifically, we propose to focus on players’ Intentions Malle et al. (2001) at various phases of the game, since comprehension of one’s and others’ intentions plays a critical role in games Goodie et al. (2012). While previous research has explored intention detection (Zhang et al., 2020; Casanueva et al., 2020), the field remains largely underexplored in game contexts. As shown in Figure 1, we design four tests: (1) Intention Selection, (2) Intention Following, (3) Intention Summarization, (4) Intention Guessing, respectively for the four social intelligence dimensions.
@@ 2.1 LLM-Based Avalon Game
Avalon is a social deduction game where players are either loyal or evil, aiming to succeed in or sabotage quests through strategy, persuasion, deduction, and deception. The primary goal is for loyal players to complete quests, while evil players aim to fail them. [^1] The detailed rules of Avalon can be found in https://avalon.fun/pdfs/rules.pdf. Leveraging the publicly available prompts proposed by Wang et al. (2023), which includes first-order and second-order reasoning before speaking, we build a framework allowing five to ten players in gameplay (we set the players number as five in this paper which includes 2 evil players and 3 loyal players). As shown in Figure 2, each game round consists of five principal components: summarizing previous rounds, discussing strategies, reconsidering the team proposal by the leader, voting, and executing the quest. Notably, we introduce an additional component, the leader’s reconsideration phase as a critical, previously unexplored component in Avalon literature. [^2] The code will be released upon publication.
@@ 3.1 Situational Awareness: Intention Selection
Situational awareness is the ability to perceive environmental elements over time and space, understand their significance, and predict their future status Li et al. (2024); Endsley (1995). In our study, we assess the situational awareness of LLMs in social contexts by evaluating their intention selection. LLMs are requested to select intentions based on summarization, first-order reasoning, and the ongoing dialogue in the current round. Ideally, LLMs should exhibit a sharp awareness of dynamic interactions, choosing intentions that are contextually appropriate rather than contradictory. We deem intentions unreasonable if they are inconsistent with:
@@
We use binary indicators to evaluate the reasonableness of intentions, assigning 1 to reasonable intentions and 0 to unreasonable ones.
@@ 4.1 Human Annotation
We recruit five computer science master’s students for annotation, each assigned seven to eight game contexts. [^3] Hourly payment for all annotators in this work is \$16. Unimpactful intentions within each context are masked to maintain focus. To calculate inter-rater reliability, two files are commonly assigned across all annotators. The Fleiss’ Kappa Fleiss (1971) is calculated pairwise, with averages derived across all pairs. For the intention following, we group the scores into two categories: 1$\sim$2 and 3$\sim$5. The scores are shown in Table 3.
@@ 4.3 Results: Intention Summarization and Guessing
Table 4: Self-awareness results. The results are GPT-3.5 and GPT-4 summarize their own intentions respectively. We also show the mean and standard deviation of human study results since we include 3 human annotators.
@@ 5.1 Social Intelligence in LLMs
Artificial social intelligence has long been a significant topic (Bainbridge et al., 1994; Kovač et al., 2023; Grossmann et al., 2023; McDonald and Pearson, 2019) as it enables machines to understand and respond to human emotions and social cues, enhancing human-computer interaction and facilitating social science study (Ziems et al., 2024; Griffin et al., 2023; Dubova, 2022; Gweon et al., 2023). Gweon et al. (2023) proposed a dataset to evaluate social reasoning and Nematzadeh et al. (2018) crafted a dataset following Sally-Anne experiment to access ToM ability. Research has demonstrated the emergence and limitations of ToM abilities in LLMs (Kosinski, 2023; Moghaddam and Honey, 2023; Sap et al., 2022; Wu et al., 2023). Although LLMs can understand human intent and choose correct actions in some tasks, their overall performance remains inferior to that of humans.
@@ 5.3 Social Game Context
With the enhancement of LLM performance, more research shifts to social games that necessitate conversation and cooperation (Ibraheem et al., 2022; Mansbach et al., 2021; Meta et al., 2022; O’Gara, 2023; Park et al., 2023). Zhou et al. 2023b used reinforcement learning to incorporate intent and ToM in DnD games. Recent works focus on the two most popular board games Avalon Lan et al. (2023); Shi et al. (2023); Wang et al. (2023); Light et al. (2023b) and Werewolf Xu et al. (2023b, a), exploring social behaviors and strategic playing ability of conversational agents. Unlike previous work which aimed to improve the game performance, we focus on the in-depth social intelligence evaluation utilizing the Avalon game.
@@ 6 Conclusion
This study includes several limitations. First, we evaluate the four essential components of social intelligence but leave several aspects unexplored. For instance, the capability of a model to modify its intentions post-reflection, i.e., the Self-Correction, could also be assessed using our framework. Future research should consider expanding this approach to include more nuanced facets of social intelligence. The second limitation lies in our annotation criteria, which are restricted to assessing the reasonableness of intentions to avoid subjectivity. However, dimensions such as creativity and ingenuity are equally important. Future studies should aim to measure these attributes to ensure that models not only perform tasks effectively but also demonstrate innovative behaviors indicative of genuine intelligence. The third limitation stems from the cost. Half of our evaluations rely on costly human annotation, and the use of GPT-4 incurs significant resource demands due to the extensive context involved. To make the evaluation process more efficient and cost-effective, we need to develop methods that allow simpler tasks to be managed by smaller models, which introduces additional challenges.
@@ Appendix A Models and Paramters
We only test GPT-3.5 and GPT-4 in our work as those two are the most common models in the social game study. We also tried open-source model like LLaMA-2, however, LLaMA-2 could not follow the format instructions properly, as also observed in Wang et al. (2023). We set the temperature as 0.8 for both GPT-3.5 and GPT-4. For intention summarization, we also tried sampling five responses and got the aggregated result by majority vote. The result on multiple responses is only 1% higher than the result on single responses. Therefore, we only generate a single response for self-awareness and ToM tasks. The cost of GPT-3.5 is approximate \$80 and \$25 for GPT-4.
@@ B.1 Creation of Intention Set
The intentions were created as a combination of domain research and prompting GPT-3.5 to select intention categories from the options listed in Table 9 and generate intentions based on the selected categories. For intention category selection we use the prompt shown in Table 7 and for intention generation we use the prompt shown in Table 8. The final set of intentions is listed in Table 9.
@@ Appendix C Game Performance
Results in Table 6 indicate that loyal players are at a disadvantage with a notably low game win rate, which corroborates findings from Wang et al. (2023). While the quest win and engagement rates show a balanced performance between loyal and evil players, the evil side demonstrates superior accuracy in team selection. This observation is consistent with their strategic role in the game. The high rate of team proposal changes across all settings suggests a dynamic framework where players consider the discussions of others, an essential aspect of social games. Additionally, the high rate of Merlin assassinations significantly boosts the win rate for evil players. Our objective is not to enhance game performance per se, but to demonstrate that with an intention-guided mechanism, game performance is similar to existing studies, providing a robust testbed for studies in social intelligence.
@@ Appendix F Human Annotation Details
First answer the following questions: 1. What is your main goal(main intent) for the game? 2. What have you observed from discussions in this round? 3. What should be the ideal team composition for your side to win? 4. After analyzing the responses to previous questions, what do you need to do to win from here? Consider answers to your previous questions and your roles and characteristics and generate intentions that help you achieve your main goal (main intention). Please do not repeat intentions and ensure they are different from each other. Each intention should be 10-15 words long. Let’s think step by step before making your decisions. Remember to generate at least 3 intentions, minimum one intention for each intention category you have selected below: [[[[intentioncategories selected by player]]]]

### v3

@@ Abstract
Large language models (LLMs) have demonstrated the potential to mimic human social intelligence. However, most studies focus on simplistic and static self-report or performance-based tests, which limits the depth and validity of the analysis. In this paper, we developed a novel framework, InterIntent, to assess LLMs’ social intelligence by mapping their ability to understand and manage intentions in a game setting. We focus on four dimensions of social intelligence: situational awareness, self-regulation, self-awareness, and theory of mind. Each dimension is linked to a specific game task: intention selection, intention following, intention summarization, and intention guessing. Our findings indicate that while LLMs exhibit high proficiency in selecting intentions, achieving an accuracy of 88%, their ability to infer the intentions of others is significantly weaker, trailing human performance by 20%. Additionally, game performance correlates with intention understanding, highlighting the importance of the four components towards success in this game. These findings underline the crucial role of intention understanding in evaluating LLMs’ social intelligence and highlight the potential of using social deduction games as a complex testbed to enhance LLM evaluation. InterIntent contributes a structured approach to bridging the evaluation gap in social intelligence within multiplayer games. [^1] Code is available at https://github.com/uscnlp-lime/Inter-Intent
@@ 1 Introduction
Despite adaptations for interactive social environments Zhou et al. (2023c, a); Park et al. (2023), most existing evaluations on LLMs Le et al. (2019); Shapira et al. (2023) focus on straightforward and static daily scenarios without explicitly defined goals for them. To enhance the contexts with complexity and dynamics with clear goals for LLMs, we turn our attention to social deduction games. These games require social interactions among players, providing a more diverse test bed for LLM evaluation Qiao et al. (2023). Existing studies have successfully utilized various games to analyze LLMs’ social behaviors, such as deception and confrontation Liang et al. (2023); Ibraheem et al. (2022); Mansbach et al. (2021); Meta et al. (2022); O’Gara (2023); Xu et al. (2023a, b); Wu et al. (2024a). However, these studies often focus on ad-hoc post-analysis of game performance and overlook systematic measurement of social intelligence, limiting the comprehensive understanding of LLMs’ capabilities in social environments.
In this paper, we leverage one of the representatives in social deduction games, Avalon Light et al. (2023a); Wang et al. (2023), as the context for LLM evaluation. Avalon is a social deduction game that relies on conversation, making it an ideal test bed with its goal-driven objectives and complex mechanisms, facilitating interactions among LLMs in a grounded environment. We systematically evaluate LLMs’ social intelligence across the four aforementioned dimensions. Specifically, we propose to focus on players’ Intentions Malle et al. (2001) at various phases of the game, since comprehension of one’s and others’ intentions plays a critical role in games Goodie et al. (2012). While previous research has explored intention detection (Zhang et al., 2020; Casanueva et al., 2020), the field remains largely underexplored in game contexts. As shown in Figure 1, we design four tests: (1) Intention Selection, (2) Intention Following, (3) Intention Summarization, (4) Intention Guessing, respectively for the four social intelligence dimensions.
@@ 2.1 LLM-Based Avalon Game
Avalon is a social deduction game where players are either loyal or evil, aiming to succeed in or sabotage quests through strategy, persuasion, deduction, and deception. The primary goal is for loyal players to complete quests, while evil players aim to fail them. [^2] The detailed rules of Avalon can be found in https://avalon.fun/pdfs/rules.pdf. Leveraging the publicly available prompts proposed by Wang et al. (2023), which includes first-order and second-order reasoning before speaking, we build a framework allowing five to ten players in gameplay (we set the players number as five in this paper which includes 2 evil players and 3 loyal players). As shown in Figure 2, each game round consists of five principal components: summarizing previous rounds, discussing strategies, reconsidering the team proposal by the leader, voting, and executing the quest. Notably, we introduce an additional component, the leader’s reconsideration phase as a critical, previously unexplored component in Avalon literature.
@@ 3.1 Situational Awareness: Intention Selection
Situational awareness is the ability to perceive environmental elements over time and space, understand their significance, and predict their future status Li et al. (2024a); Endsley (1995). In our study, we assess the situational awareness of LLMs in social contexts by evaluating their intention selection. LLMs are requested to select intentions based on summarization, first-order reasoning, and the ongoing dialogue in the current round. Ideally, LLMs should exhibit a sharp awareness of dynamic interactions, choosing intentions that are contextually appropriate rather than contradictory. We deem intentions unreasonable if they are inconsistent with:
@@
We use binary indicators to evaluate the reasonableness of intentions, assigning 1 to reasonable intentions and 0 to unreasonable ones. Previous hallucinations, such as manipulated information, might influence the current player’s decisions. However, we maintain strict criteria, as we expect the player to be capable of recognizing these hallucinations.
@@ 4.1 Human Annotation
We recruit five computer science master’s students for annotation, each assigned seven to eight game contexts. [^3] Hourly payment for all annotators in this work is \$16. Unimpactful intentions within each context are masked to maintain focus. To calculate inter-rater reliability, two files are commonly assigned across all annotators. The Fleiss’ Kappa Fleiss (1971) is calculated pairwise, with averages derived across all pairs. For the intention following, we group the scores into two categories: 1$\sim$3 and 4$\sim$5. Scores of 5 represent a perfect response, while 4 indicates a nearly perfect response. Scores from 1 to 3 signify that the response contains significant flaws. The scores are shown in Table 3. Additional grouping results are provided in the Appendix Sec F.
@@ 4.3 Results: Intention Summarization and Guessing
Table 4: Self-awareness results. The results are GPT-3.5 and GPT-4 summarize their own intentions respectively. We also show the mean and standard deviation of human study results since we include 3 human annotators. We show $F_{1}$ score in this table.
@@ 5.1 Social Intelligence in LLMs
Artificial social intelligence has long been a significant topic (Bainbridge et al., 1994; Kovač et al., 2023; Grossmann et al., 2023; McDonald and Pearson, 2019; Street, 2024; Li et al., 2024b) as it enables machines to understand and respond to human emotions and social cues, enhancing human-computer interaction and facilitating social science study (Ziems et al., 2024; Griffin et al., 2023; Dubova, 2022; Gweon et al., 2023; Zhang et al., 2024). Gweon et al. (2023) proposed a dataset to evaluate social reasoning and Nematzadeh et al. (2018) crafted a dataset following Sally-Anne experiment to access ToM ability. Research has demonstrated the emergence and limitations of ToM abilities in LLMs (Kosinski, 2023; Moghaddam and Honey, 2023; Sap et al., 2022; Wu et al., 2023). Although LLMs can understand human intent and choose correct actions in some tasks, their overall performance remains inferior to that of humans.
@@ 5.3 Social Game Context
With the enhancement of LLM performance, more research shifts to social games that necessitate conversation and cooperation (Ibraheem et al., 2022; Mansbach et al., 2021; Meta et al., 2022; O’Gara, 2023; Park et al., 2023; Zhou et al., 2024; Wu et al., 2024b). Zhou et al. 2023b used reinforcement learning to incorporate intent and ToM in DnD games. Recent works focus on the two most popular board games Avalon Lan et al. (2023); Shi et al. (2023); Wang et al. (2023); Light et al. (2023b) and Werewolf Xu et al. (2023b, a), exploring social behaviors and strategic playing ability of conversational agents. Unlike previous work which aimed to improve the game performance, we focus on the in-depth social intelligence evaluation utilizing the Avalon game.
@@ 6 Conclusion
This study includes several limitations. First, we evaluate the four essential components of social intelligence but leave several aspects unexplored. For instance, the capability of a model to modify its intentions post-reflection, i.e., the Self-Correction, could also be assessed using our framework. Future research should consider expanding this approach to include more nuanced facets of social intelligence. The second limitation lies in our annotation criteria, which are restricted to assessing the reasonableness of intentions to avoid subjectivity. However, dimensions such as creativity and ingenuity are equally important. Future studies should aim to measure these attributes to ensure that models not only perform tasks effectively but also demonstrate innovative behaviors indicative of genuine intelligence. The third limitation stems from the cost. Half of our evaluations rely on costly human annotation, and the use of GPT-4 incurs significant resource demands due to the extensive context involved. To make the evaluation process more efficient and cost-effective, we need to develop methods that allow simpler tasks to be managed by smaller models, which introduces additional challenges. The final limitation is that we did not use an open-source model, as it lacks the capability to effectively engage in social games.
@@ Appendix A Models and Paramters
We only test GPT-3.5 and GPT-4 in our work as those two are the most common models in the social game study. We also tried open-source model like LLaMA-2 7B and LLaMA-3.1 8B, however, LLaMA-2 could not follow the format instructions properly, as also observed in Wang et al. (2023). LLaMA-3.1 performs better than LLaMA-2 but still fails to follow the format after a few turns of conversations. We set the temperature as 0.8 for both GPT-3.5 and GPT-4. For intention summarization, we also tried sampling five responses and got the aggregated result by majority vote. The result on multiple responses is only 1% higher than the result on single responses. Therefore, we only generate a single response for self-awareness and ToM tasks. The cost of GPT-3.5 is approximately \$80 and \$25 for GPT-4.
@@ B.1 Creation of Intention Set
The intentions were created as a combination of domain research and prompting GPT-3.5 to select intention categories from the options listed in Table 9 and generate intentions based on the selected categories. For intention category selection we use the prompt shown in Table 7 and for intention generation, we use the prompt shown in Table 8. The final set of intentions is listed in Table 9.
@@ Appendix C Game Performance
We record the entire game and extract intermediate steps, such as voting results, team proposals, and quest outcomes, to calculate game performance. Results in Table 6 indicate that loyal players are at a disadvantage with a notably low game win rate, which corroborates findings from Wang et al. (2023). While the quest win and engagement rates show a balanced performance between loyal and evil players, the evil side demonstrates superior accuracy in team selection. This observation is consistent with their strategic role in the game. The high rate of team proposal changes across all settings suggests a dynamic framework where players consider the discussions of others, an essential aspect of social games. Additionally, the high rate of Merlin assassinations significantly boosts the win rate for evil players. Our objective is not to enhance game performance per se, but to demonstrate that with an intention-guided mechanism, game performance is similar to existing studies, providing a robust testbed for studies in social intelligence.
@@ Appendix F Human Annotation Details
We experimented with several grouping methods to calculate agreement scores. When using the grouping of 1 $\sim$ 4 and 5, the Fleiss’ Kappa scores for "thinking" and "speaking" were 0.448 and 0.420, respectively. These scores are relatively lower, likely due to differing interpretations of informativeness among human annotators. For the grouping of 1 $\sim$ 2 and 3 $\sim$ 5, the agreement was low, as scores of 1 $\sim$ 2 accounted for only a small portion of the entire dataset. Since Fleiss’ Kappa is highly sensitive to disagreements within the 1 $\sim$ 2 group, even minor discrepancies resulted in an overall lower score.
@@
First answer the following questions: 1. What is your main goal(main intent) for the game? 2. What have you observed from discussions in this round? 3. What should be the ideal team composition for your side to win? 4. After analyzing the responses to previous questions, what do you need to do to win from here? Consider answers to your previous questions and your roles and characteristics and generate intentions that help you achieve your main goal (main intention). Please do not repeat intentions and ensure they are different from each other. Each intention should be 10-15 words long. Let’s think step by step before making your decisions. Remember to generate at least 3 intentions, with a minimum of one intention for each intention category you have selected below: [[[[intentioncategories selected by player]]]]